# Evaluating Large Language Models on Time Series Feature Understanding: A Comprehensive Taxonomy and Benchmark

### v1

@@ 6.2 Benchmark Results
In Table 2, we display the main results for all tasks outlined in Sec. 4.
Metric GPT4 GPT3.5 Llama2 Vicuna
Univariate time series characteristics
Feature detection
Trend F1score 0.88 0.43 0.54 0.60
Seasonality F1score 0.98 0.70 0.71 0.47
Outlier F1score 0.53 0.53 0.46 0.53
Struct. break F1score 0.67 0.56 0.43 0.52
Volatility F1score 0.45 0.50 0.45 0.50
Fat Tails F1score 0.43 0.51 0.31 0.44
Stationarity F1score 0.31 0.31 0.31 0.31
Feature classification
Trend F1score 0.98 0.47 0.41 0.61
Seasonality F1score 0.25 0.15 0.17 0.20
Outlier F1score 0.67 0.17 0.07 0.28
Struct. break F1score 0.34 0.48 0.31 0.36
Volatility F1score 0.13 0.16 0.10 0.23
Multivariate time series characteristics
Fixed Corr. F1score 0.40 0.42 0.30 0.32
Lagged Corr. F1score 0.44 0.47 0.22 0.33
Changing Corr. F1score 0.43 0.41 0.23 0.41
Information Retrieval
Value on Date Acc 1.00 0.94 0.38 0.48
Value on Date MAPE 0.00 0.10 0.65 0.78
Arithmetic Reasoning
Min Value Acc 1.00 0.99 0.58 0.66
MAPE 0.00 0.04 16.18 12.24
Min Date Acc 0.98 0.94 0.38 0.55
Max Value Acc 0.97 0.92 0.56 0.46
MAPE 0.01 0.08 0.95 0.74
Max Date Acc 0.96 0.88 0.46 0.42
@@
The results for univariate time series feature detection and classification tasks illustrate GPT4’s robustness in trend and seasonality detection, substantially outperforming Llama2 and Vicuna. However, the detection of structural breaks and volatility presents challenges across all models, with lower accuracy scores.
GPT4 excels in trend classification tasks, demonstrating superior performance. However, in classifying seasonality, outliers, and structural breaks, performance is mixed, with Vicuna sometimes surpassing Llama2, highlighting the distinct strengths of each model. Figure 4 summarizes the accuracy performance for the information retrieval and arithmetic reasoning tasks, and F1 score for the feature detection and classification tasks for all models.
@@
In multivariate time series feature detection and classification tasks, all models achieve moderate accuracy, suggesting potential for enhancement in intricate multivariate data analysis.
@@
For information retrieval tasks, GPT4 outperforms GPT3.5 and other models, achieving perfect accuracy in identifying the value on a given date. It also maintains a low Mean Absolute Percentage Error (MAPE), indicative of its precise value predictions. The arithmetic reasoning results echo these findings, with GPT4 displaying superior accuracy, especially in determining minimum and maximum values within a series.
(a) IR and arithmetic reasoning
(b) Feature detection
(c) Feature classification
@@
Figure 4: Feature detection and classification scores of GPT4, GPT3.5, Vicuna and Llama2.
@@ 6.3 Deep Dive on Performance Factors
We present four formatting approaches in this section, csv, which is a common comma separated value, plain where the time series is formatted as Date:YYYY-MM-DD,Value:num for each pair date-value. We also use the formatting approach proposed by Gruver et al. (2023) which we denominate spaces that adds blank spaces between each digit of the time series, tokenizing each digit individually, and symbol, an enriched format where we add a column to the time series with arrows indicating if the value has moved up, down or remained unchanged. Examples of every approach can be found in Sec. E in the Appendix.
Table 3 shows the results for the four time series formatting strategies. For the full results, please refer to Tables 10 and 9. For the information retrieval and arithmetic reasoning tasks, the plain formatting yields better results across all models. This approach provides more structure to the input, and outperforms other formats in a task where the connection between time and value is important. For the detection and classification tasks, the plain formatting does not yield better results. Interestingly the symbol formatting that adds an additional column to the time series yields better results in the trend classification task. This means that the LLMs can correctly map the symbol to the time series movement and use it to achieve the best performance in trend classification. Furthermore, GPT3.5 leverages this additional information in the trend and anomalies datasets but not in the seasonality dataset.
@@
Time Series Length
Figure 5 shows the performance of GPT3.5, Llama2 and Vicuna on three datasets, trend, seasonality and outliers wich have time series with different lengths. We observe that GPT3.5 retrieval performance degrades slowly with increasing sequence length. Llama2 and and Vicuna suffer a more steep degradation especially from time series of length 30 steps to 60 steps; for longer sequences the degradation in performance becomes linear.
@@
We carry out a series of experiments to determine how the position of the target value affects task performance across various types of time series data. We address progressively more complex objectives: 1) identifying the presence of a value in a time series without a specified date (D.1); 2) retrieving a value corresponding to a specific date (D.2); and 3) identifying the minimum and maximum values (D.3). We cover a range of time series data, from monotonic series without noise to those with noise, sinusoidal patterns, data featuring outliers (spikes), and Brownian motion scenarios, each adding a layer of complexity. We examine how the position of the target value within the four quadrants — 1st, 2nd, 3rd, and 4th— affects the efficacy of these tasks across the varied time series landscapes. This approach helps reveal the influence of position on different LLMs (GPT3.5, Llama2, and Vicuna) in the task of time series understanding.
We consider the presence of position bias when the maximum performance gap between quadrants exceeds 10%. Given this criterion, our analysis provides the following key takeaways on position bias impacting LLM performance across the defined tasks:
  • Pronounced position bias is observed across all tasks and LLMs
    • GPT models show significant bias exclusively in complex tasks that involve arithmetic reasoning.
    • Both Llama2 and Vicuna demonstrate position biases across all tasks, from the simplest to the most complex ones.
  • The degree of complexity in the time series data tends to increase the extent of position bias observed within each task.
Refer to Section D in the appendix, where we offer a detailed analysis of position bias across each task to further substantiate these conclusions.
@@ 7 Conclusion
In conclusion, we provide a critical examination of general-purpose Large Language Models (LLMs) in the context of time series understanding. Through the development of a comprehensive taxonomy of time series features and the synthesis of a diverse dataset that encapsulates these features, we have laid a solid foundation for evaluating the capabilities of LLMs in understanding and interpreting time series data. Our systematic evaluation sheds light on the inherent strengths and limitations of these models, offering valuable insights for practitioners aiming to leverage LLMs in time series understanding. Recognizing the areas of weakness and strength in general-purpose LLMs’ current capabilities allows for targeted enhancements, ensuring that these powerful models can be more effectively adapted to specific domains.
@@ Acknowledgements
Appendix A Synthetic Time Series Dataset
A.1 Univariate Time Series
@@
    Volatility We generated synthetic time series data to simulate various volatility patterns, specifically targeting clustered volatility, leverage effects, constant volatility, and increasing volatility, to mimic characteristics observed in financial markets. For clustered volatility, we utilized a GARCH(1,1) model with parameters $\omega=0.1$, $\alpha=0.05$, and $\beta=0.9$, ensuring the sum of $\alpha$ and $\beta$ remained just below 1 for stationarity, thus capturing high volatility persistence. To simulate the leverage effect, our model increased volatility in response to negative returns, reflecting typical market dynamics. Additionally, we created time series with constant volatility by adding normally distributed random noise (standard deviation of 1) to a cumulative sum of random values, and with increasing volatility by scaling the noise in proportion to the increasing range of the series (scaling factor up to 5 towards the end of the series). The latter was achieved by multiplying the standard deviation of the random noise by a linearly increasing factor, resulting in a volatility profile that progressively intensified. These methodologies enabled us to comprehensively represent different volatility behaviors in financial time series, including constant, increasing, clustered, and leverage-induced volatilities, thereby enriching our analysis with diverse market conditions.
@@
A.2 Multivariate Time Series
@@ A.3 Data Examples
[Uncaptioned image] (a) Linear trend [Uncaptioned image] (b) Quadratic trend [Uncaptioned image] (c) Exponential trend [Uncaptioned image] (d) No clear trend
@@
[Uncaptioned image] (a) Fixed seasonality [Uncaptioned image] (b) Varying amplitude [Uncaptioned image] (c) Shifting patterns [Uncaptioned image] (d) Multiple Seasonalities
@@
Appendix B Additional datasets
@@
Appendix C Additional results
C.1 Trend
@@ C.4 Volatility
Appendix D Position Bias
D.1 Does the position of the target value affect the performance of identifying its presence in various types of time series data?
@@ Appendix D Position Bias
D.2 Does the position impact the retrieval performance for a specific date’s value from time series data?
@@
Once again, GPT achieves nearly perfect performance across all quadrants and time series types, suggesting no position bias in the retrieval task either. Similar to the findings in D.1, Vicuna outperforms Llama2. Moreover, both Vicuna and Llama2 exhibit position bias in most datasets, including monotonic series both with and without noise, and sinusoidal series.
@@
D.3 Does the position impact the efficiency of identifying minimum and maximum values in different types of time series data?
@@
Appendix E Time Series formatting

### v2

@@ 6.2 Benchmark Results
In Table 2, we display the main results for the feature detection, feature classification, information retrieval and arithmetic reasoning tasks outlined in Sec. 4. The results for univariate time series feature detection and classification tasks illustrate GPT4’s robustness in trend and seasonality detection, substantially outperforming Llama2, Vicuna, and GPT3.5 in zero-shot settings. This performance is further enhanced when chain-of-thought prompting is used. However, the detection of structural breaks and volatility presents challenges across all models, with lower accuracy scores, even with chain-of-thought prompting. GPT4 tends to always answer no for stationarity and fat tail detection tasks, while in the case of chain-of-thought prompting it does not answer, clarifying that it is only an AI model and cannot perform the necessary statistical tests.
For trend classification, GPT4 excels in zero-shot and chain-of-thought prompting, demonstrating superior performance. Phi3 shows strong performance in zero-shot settings for trend classification, even surpassing GPT3.5 in zero-shot. In classifying seasonality, outliers, and structural breaks, Phi3 also demonstrates competitive performance, sometimes surpassing Llama2 and Vicuna, and outperforming GPT3.5 in seasonality classification, highlighting its distinct strengths. Additional plots of confusion matrices are provided in Appendix D to better understand how the models select their choices, revealing potential biases such as consistently selecting the same label. Figure 2 (a) summarizes the F1 score for the feature detection task for all models, showing the strong performance on the four easier features, with Phi3 also being competitive in trend, seasonality and volatility detection.
@@
In multivariate time series feature detection and classification tasks, all models achieve moderate accuracy in zero-shot settings, suggesting potential for enhancement in intricate multivariate data analysis. Chain-of-thought prompting does not significantly improve performance in this context.
@@
For information retrieval tasks, GPT4 outperforms GPT3.5 and other models, achieving perfect accuracy in identifying the value on a given date. It also maintains a low Mean Absolute Percentage Error (MAPE), indicative of its precise value predictions. The arithmetic reasoning results echo these findings, with GPT4 displaying superior accuracy, especially in determining minimum and maximum values within a series. Figure 2 summarizes the accuracy performance for the information retrieval and arithmetic reasoning tasks, where there are two clear groups with similar performance, GPT4, GPT3.5 and Phi3, and Llama2 and Vicuna.
(a) Feature detection
(b) IR and math reasoning
@@
Figure 2: Feature detection and arithmetic reasoning scores of GPT4, GPT3.5, Vicuna, Llama2 and Phi3.
In the text matching tasks, Table 3(a) shows results intra-datasets, where GPT-4 significantly outperforms other models, achieving near-perfect accuracy across all datasets. This suggests that GPT-4 is capable of understanding the nuances of both qualitative and quantitative time series descriptions and effectively relating them to the underlying data. Table 3(b) shows the results for the matching cross-datasets where GPT-4 outperforms other models on all datasets except two, showcasing its superior capability in understanding and matching qualitative descriptions even without explicit quantitative cues. The performance of GPT-3.5, Llama2, Vicuna, and Phi-3 is notably lower, indicating a greater reliance on quantitative information for accurate matching in these models. This overall decrease in performance, is in line with our overall findings that while numerical performance on simple arithmetic tasks is quite high, performance is generally lower for time series feature detection and classification.
Table 3: Accuracy of LLMs in matching time series to their corresponding textual descriptions, given four options. (Bold indicates best performance)
GPT-4 GPT-3.5 Llama2 Vicuna Phi3
Trend 1.00 0.74 0.67 0.53 0.73
Seasonality 0.93 0.64 0.58 0.47 0.64
Anomalies 1.00 0.69 0.62 0.47 0.69
Struct. break 0.99 0.63 0.57 0.39 0.63
Volatility 0.98 0.72 0.60 0.49 0.65
Stationarity 0.99 0.72 0.64 0.52 0.69
Fat Tails 0.99 0.69 0.61 0.43 0.68
(a) Intra-dataset matching
GPT-4 GPT-3.5 Llama2 Vicuna Phi3
Trend 0.46 0.21 0.32 0.36 0.34
Seasonality 0.41 0.50 0.32 0.35 0.31
Anomalies 0.46 0.16 0.32 0.36 0.34
Struct. break 0.28 0.1 0.26 0.27 0.24
Volatility 0.10 0.07 0.15 0.12 0.14
Stationarity 0.53 0.53 0.36 0.42 0.35
Fat Tails 0.10 0.04 0.10 0.10 0.09
(b) Cross-dataset matching
@@ 6.3 Deep Dive on Performance Factors
We present four formatting approaches in this section, csv, which is a common comma separated value, plain where the time series is formatted as Date:YYYY-MM-DD,Value:num for each pair date-value. We also use the formatting approach proposed by Gruver et al. (2023) which we denominate spaces that adds blank spaces between each digit of the time series, tokenizing each digit individually, and symbol, an enriched format where we add a column to the time series with arrows indicating if the value has moved up, down or remained unchanged. Examples of every approach can be found in Sec. F in the Appendix.
Table 4 shows the results for the four time series formatting strategies. For the information retrieval and arithmetic reasoning tasks, the plain formatting yields better results across all models. This approach provides more structure to the input, and outperforms other formats in a task where the connection between time and value is important. For the detection and classification tasks, the plain formatting does not yield better results. Interestingly the symbol formatting that adds an additional column to the time series yields better results in the trend classification task. This indicates that LLMs can effectively leverage symbolic representations of time series movements to enhance their understanding in trend classification.
@@
Time Series Length
Figure 3 shows the performance of GPT3.5, Phi3, Llama2 and Vicuna on three datasets, trend, seasonality and outliers which have time series with different lengths. We observe that GPT3.5 and Phi3 retrieval performance degrades slowly with increasing sequence length. Llama2 and and Vicuna suffer a more steep degradation especially from time series of length 30 steps to 60 steps.
@@
We carry out a series of experiments to determine how the position of the target value affects task performance across various types of time series data. We address progressively more complex objectives: 1) identifying the presence of a value in a time series without a specified date (E.1); 2) retrieving a value corresponding to a specific date (E.2); and 3) identifying the minimum and maximum values (E.3). We cover a range of time series data, from monotonic series without noise to those with noise, sinusoidal patterns, data featuring outliers (spikes), and Brownian motion scenarios, each adding a layer of complexity. We examine how the position of the target value within the four quadrants — 1st, 2nd, 3rd, and 4th— affects the efficacy of these tasks across the varied time series landscapes. This approach helps reveal the influence of position on different LLMs (GPT3.5, Llama2, and Vicuna) in the task of time series understanding.
We consider the presence of position bias when the maximum performance gap between quadrants exceeds 10%. Given this criterion, our analysis provides the following key takeaways on position bias impacting LLM performance across the defined tasks: (1) Pronounced position bias is observed across all tasks and LLMs: GPT models show significant bias exclusively in complex tasks that involve arithmetic reasoning. Both Llama2 and Vicuna demonstrate position biases across all tasks, from the simplest to the most complex ones. (2) The degree of complexity in the time series data tends to increase the extent of position bias observed within each task. See Appendix E, where we offer a detailed analysis of position bias across each task to further substantiate these conclusions.
@@ 7 Conclusion
In conclusion, we provide a critical examination of general-purpose Large Language Models (LLMs) in the context of time series understanding. Through the development of a comprehensive taxonomy of time series features and the synthesis of a diverse dataset that encapsulates these features, including qualitative and quantitative textual descriptions for each time series, we have laid a solid foundation for evaluating the capabilities of LLMs in understanding and interpreting time series data. Our systematic evaluation sheds light on the inherent strengths and limitations of these models, offering valuable insights for practitioners aiming to leverage LLMs in time series understanding. Recognizing the areas of weakness and strength in general-purpose LLMs’ current capabilities allows for targeted enhancements, ensuring that these powerful models can be more effectively adapted to specific domains.
In the future, we plan to study the performance of LLMs on real-world time series datasets to assess the generalizability of the proposed framework. This will involve testing LLMs on diverse datasets from various domains, such as finance, healthcare, and climate science. Additionally, future work should expand the analysis challenges LLMs face with multivariate time series data, including the ability to identify and interpret relationships between multiple series, such as correlation, cross-correlation, and dynamic conditional correlation. Understanding these challenges will be crucial for developing more effective LLMs for complex time series analysis. Finally, evaluating LLMs in few-shot settings is an important area for future work, as it can reveal the models’ ability to learn and generalize from limited time series data. This can be particularly valuable in domains where labeled data is scarce or expensive to obtain.
@@ Acknowledgements
Appendix A Additional details of Taxonomy
Feature Description Example Use Cases
Trend The general direction of a time series, either increasing (upward) or decreasing (downward) over a long period. Finance: Stock price trends, inflation rates. Climate: Global temperature trends. Energy: Long-term energy consumption trends.
Seasonality A repeating pattern in a time series that occurs at regular intervals, such as daily, weekly, monthly, or yearly. Energy: Seasonal variations in electricity demand. Retail: Seasonal sales patterns (e.g., holiday shopping). Tourism: Seasonal fluctuations in visitor numbers.
Fixed-Period Seasonality with a constant, unchanging period (e.g., monthly seasonality). Energy: Monthly variations in electricity usage. Finance: Quarterly earnings reports.
Shifting Period Seasonal patterns where the length of the period shifts over time. Climate: Shifting seasonal temperature patterns due to climate change. Retail: Shifting sales patterns due to changing consumer behavior.
Multiple Seasonality Presence of multiple overlapping seasonal patterns (e.g., both weekly and monthly seasonality). Finance: Weekly and monthly trading cycles. Health: Weekly and annual cycles in flu cases.
Volatility The degree of variation of a time series over time, often measured by the standard deviation or variance. Finance: Stock market volatility, exchange rate fluctuations. Energy: Price volatility in commodity markets. Weather: Day-to-day fluctuations in temperature or precipitation.
Constant Volatility The degree of variation in the time series remains consistent and predictable over time. Finance: Stable bond markets. Energy: Consistent electricity prices.
Trending Volatility The level of variation in the time series shows a clear increasing or decreasing trend over time. Finance: Increasing volatility in emerging markets. Climate: Increasing variability in weather patterns.
Clustered Volatility The time series exhibits periods where volatility is significantly higher or lower, with these periods tending to cluster together. Finance: Volatility clustering in financial markets during crises. Economics: Clustered periods of high inflation.
Dynamic Volatility The volatility of the time series changes over time in response to external factors (e.g., leverage effect where the volatility of the time series tends to increase when the series experiences negative returns). Finance: Changing volatility due to market interventions. Climate: Volatility changes in response to natural disasters.
Anomalies Data points that deviate significantly from the expected pattern of a time series. Quality Control: Detecting defective products in a manufacturing process. Network Security: Identifying unusual traffic patterns that may indicate cyberattacks. Finance: Detecting fraudulent transactions.
Spike A sudden and brief deviation from the overall pattern of the data. Finance: Sudden stock price jumps. Weather: Temperature spikes during heatwaves.
Level Shift A sudden and lasting change in the average value of a time series. Economics: Changes in consumer confidence or business sentiment. Energy: Shifts in energy consumption patterns due to technological advancements or policy changes. Environmental Science: Changes in water levels or pollutant concentrations due to natural or human-induced factors.
Temporal Disruption An interval where data is missing or not recorded. Network Security: Periods of data loss in network traffic. Health: Missing data in patient records.
Structural Breaks Abrupt changes in the underlying structure of a time series, often caused by external events or policy changes. Economics: Changes in economic policy or regulations. Finance: Market crashes or financial crises. Epidemiology: Changes in disease transmission patterns due to interventions.
Stationarity A time series is stationary if its statistical properties, such as mean and variance, do not change over time. Econometrics: Assumption for many time series models. Finance: Assessing the stability of financial markets.
Fat Tails A distribution of a time series where extreme events are more likely than expected under a normal distribution. Finance: Modeling extreme price movements in financial markets. Insurance: Pricing insurance policies for catastrophic events.
Table 5: Definitions and examples of time series analysis features and sub-categories.
Appendix B Synthetic Time Series Dataset
B.1 Univariate Time Series
@@
    Volatility We generated synthetic time series data to simulate various volatility patterns, specifically targeting clustered volatility, leverage effects, constant volatility, and increasing volatility, to mimic characteristics observed in financial markets.
    For clustered volatility, we utilized a GARCH(1,1) model with parameters $\omega=0.1$, $\alpha=0.2$, and $\beta=0.7$, ensuring the sum of $\alpha$ and $\beta$ remained below 1 for stationarity, thus capturing high volatility persistence. The GARCH(1,1) model is defined by the equations:
    $\sigma_{t}^{2}=\omega+\alpha r_{t-1}^{2}+\beta\sigma_{t-1}^{2}$
    $r_{t}=\sigma_{t}\epsilon_{t}$
    where $\sigma_{t}^{2}$ is the conditional variance, $r_{t}$ is the return at time $t$, and $\epsilon_{t}$ is white noise.
    To simulate the leverage effect, our model increased volatility in response to negative returns, reflecting typical market dynamics. The leverage effect model was designed with a base volatility of 0.1 and a leverage strength of 0.3, ensuring that volatility would significantly increase after negative returns while gradually reverting to the base level after positive returns. The model is defined by:
    $r_{t}=\sigma_{t-1}\epsilon_{t}$
    $\sigma_{t}=\begin{cases}\sigma_{t-1}(1+\text{leverage\_strength})&\text{if }r_{t}<0\\ \max(\sigma_{t-1}(1-\text{leverage\_strength}),0.01)&\text{if }r_{t}\geq 0\end{cases}$
    Additionally, we created time series with constant volatility by adding normally distributed random noise (standard deviation of 1) to a cumulative sum of random values. This produced a time series with a consistent level of volatility throughout the period. Mathematically, this is represented as:
    $r_{t}=\sum_{i=1}^{t}\epsilon_{i}+\eta_{t}$
    where $\epsilon_{i}$ is white noise and $\eta_{t}\sim N(0,1)$.
    For increasing volatility, we scaled the noise in proportion to the increasing range of the series, with a scaling factor up to 5 towards the end of the series. This was achieved by multiplying the standard deviation of the random noise by a linearly increasing factor, resulting in a volatility profile that progressively intensified. This can be described by:
    $\sigma_{t}=\sigma_{0}\left(1+\frac{t}{n}\cdot 5\right)$
    $r_{t}=\epsilon_{t}\cdot\sigma_{t}$
    where $\sigma_{0}$ is the initial standard deviation and $n$ is the total number of points.
    To ensure non-negative volatility values across all simulations, we took the absolute values of the generated noise. These methodologies enabled us to comprehensively represent different volatility behaviors in financial time series, including constant, increasing, clustered, and leverage-induced volatilities. By using these varied approaches, we enriched our analysis with diverse market conditions, providing a robust dataset for evaluating the performance of models designed to handle different volatility patterns.
@@
B.2 Multivariate Time Series
@@ B.3 Data Examples
[Uncaptioned image] (a) Positive trend [Uncaptioned image] (b) Negative trend [Uncaptioned image] (c) Positive trend [Uncaptioned image] (d) No clear trend
@@
[Uncaptioned image] (a) Fixed seasonality [Uncaptioned image] (b) Fixed seasonality [Uncaptioned image] (c) Shifting patterns [Uncaptioned image] (d) Multiple Seasonalities
@@
Appendix C Additional datasets
@@
Appendix D Additional results
D.1 Trend
@@ D.4 Volatility
Appendix E Position Bias
E.1 Does the position of the target value affect the performance of identifying its presence in various types of time series data?
@@ Appendix E Position Bias
E.2 Does the position impact the retrieval performance for a specific date’s value from time series data?
@@
Once again, GPT achieves nearly perfect performance across all quadrants and time series types, suggesting no position bias in the retrieval task either. Similar to the findings in E.1, Vicuna outperforms Llama2. Moreover, both Vicuna and Llama2 exhibit position bias in most datasets, including monotonic series both with and without noise, and sinusoidal series.
@@
E.3 Does the position impact the efficiency of identifying minimum and maximum values in different types of time series data?
@@
Appendix F Time Series formatting